OpenStudy (anonymous):

if x^3=x^9, what is x in terms of y?

OpenStudy (owlcoffee):

if you want to write a variable in terms of another, it is isolating it on any side of the equality sign. Which would mean we'd have to take the thord root of both sides in order to get rid of x's exponent: \[x^3=y^9\] \[x=\sqrt[3]{y^9}\] You can leave it like that, or you can simplify further, but esscentially the answer is that.

hero (hero):

How do you know it isn't y^3 = x^9 ?

OpenStudy (owlcoffee):

Good question, I had to suppose one scenario. But the idea is the same wether it's x^3=y^9 or y^3=x^9.